OpenStudy (kanwal32):

The driver of a train is travelling at 115km/h sees on the same track,100m in front of him,a slow train travelling in the same direction at 25km/h. The least retardation that must applied to a faster train to avoid a collosion

OpenStudy (kanwal32):

@ikram002p hlp

OpenStudy (kanwal32):

@ganeshie8 hlp

OpenStudy (kanwal32):

i think differentiation is required

ganeshie8 (ganeshie8):

\[\large v^2 - u^2 = 2as\] \(u = 115-25 = 95kmph = ?m/s\) \(v=0\) \(s = 100\) plug them and solve \(a\)

OpenStudy (kanwal32):

@ganeshie8 here we will use relative motio properties

OpenStudy (kanwal32):

this will give u a wrong answer

OpenStudy (kanwal32):

sorry i didn't see ur the whone thing @ganeshie8

OpenStudy (kanwal32):

no option like this

OpenStudy (kanwal32):

@hartnn hlp

OpenStudy (kanwal32):

@nincompoop @hartnn

OpenStudy (kanwal32):

pls hlp

OpenStudy (kanwal32):

@Abhisar hlp

OpenStudy (kanwal32):

@satellite73

OpenStudy (kanwal32):

@hartnn save my life tommorrow is my exam

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @ganeshie8 \[\large v^2 - u^2 = 2as\] \(u = 115-25 =\cancel{95kmph}~90 = ?m/s\) \(v=0\) \(s = 100\) plug them and solve \(a\) \(\color{blue}{\text{End of Quote}}\)

hartnn (hartnn):

an object moving with velocity x, and another with velocity y in the same direction, then relative velocity is |x-y| an object moving with velocity x, and another with velocity y in the opposite direction, then relative velocity is x+y

OpenStudy (kanwal32):

ok its 25 i got it thnx @Abhisar @hartnn @ganeshie8

ganeshie8 (ganeshie8):

careful about the units

ganeshie8 (ganeshie8):

did u get 25 m/s^2 ?

OpenStudy (kanwal32):

no 3.125,/s^2 i was talking about velocity @ganeshie8

ganeshie8 (ganeshie8):

how did u get 3.125 m/s^2 ?

OpenStudy (kanwal32):

90*5/18 to convert in m/s and then pluging the values